# DiverseNet: Decision Diversified Semi-supervised Semantic Segmentation Networks for Remote Sensing Imagery

### v1

@@ II Cross-head Supervision (DiverseHead)
Figure 2: DiverseHead: an online semi-supervised learning approach. The slide windows of the freezer randomly select a certain number of heads to freeze the parameter update of heads. A mainstream head is randomly selected to be supervised by unsupervised loss.
@@
This section introduces the details of the proposed DiverseHead Semi-Supervised Learning method. Inspired by Bagging (also called bootstrap aggregating), DiverseHead is proposed to create various predictions by multiple heads with even the same structure. The source of perturbation is from the dynamic choosing heads to be updated for each iteration, called dynamic freezing, which is equivalent to using different training data for different heads in each iteration. The proposed DiverseHead differs from CCT, which does not distinguish the main or auxiliary decoder. Also, all decoders are trained by both supervised and unsupervised losses. In addition, the proposed method combines self-training and consistency regularization, which calculates unsupervised loss by using optimized pseudo labels. Specially, given are both a labelled dataset $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$ containing M images and an unlabelled data set $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$ with N images. The network $Q$ is constructed with multiple heads denoted as $\left\{head^{i}\right\}_{i=1}^{len(heads)}$. Each of these heads is initialized differently. The expectation is to gain the trained network $Q$ leveraging these labelled and unlabelled data based on the proposed semi-supervised learning methods.
@@
The output of each head refers to a probability-based prediction for all classifications corresponding to its input. When working with labelled data, the supervised loss $\mathcal{L}_{sup}^{j}$ between the $k\leavevmode\nobreak\ th$ ground truth $y_{k}$ and its corresponding prediction $p_{k}^{j}$ for $j\leavevmode\nobreak\ th$ head is defined by using the standard cross-entropy loss function $\ell_{ce}$:
$\mathcal{L}_{sup}^{j}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\left(\ell_{ce}\left(p_{k}^{j},y_{k}\right)\right),$ (1)
@@
where $W$ and $H$ refer to the weight and height of input images. The ultimate supervised loss is determined by the mean of the losses across all heads
$\mathcal{L}_{sup}=\frac{1}{len(heads)}\sum_{j=1}^{len(heads)}\mathcal{L}_{sup}^{j}.$ (2)
In each iteration, we also get prediction $r_{k}^{j}$ corresponding to unlabelled data $u_{k}^{j}$ from $j\leavevmode\nobreak\ th$ head. In order to get higher precision pseudo labels, inspired by ensemble learning, the predictions from all heads are aggregated to create a combined prediction $\hat{r}_{k}^{sum}$. Subsequently, this combined prediction is utilized to calculate pseudo labels $\hat{y}_{k}^{sum}$ through an $argmax$ operation, which returns the indices of the maximum values of the prediction along the class dimension. Next, there is a random selector to pick only one head as the main branch to be supervised by minimizing the unsupervised loss
@@
$\mathcal{L}_{unsup}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\left(\ell_{ce}\left(r_{k}^{main},\hat{y}_{k}^{sum}\right)\right),$ (3)
where $r_{k}^{main}$ is the prediction from the selected head.
@@
Two perturbation methods. Based on the framework of DiverseHead, we propose a parameter perturbation method called dynamic freezing and also apply dropout as another perturbation to increase the diversity of features during training. Specifically, the first step of dynamic freezing refers to a process where, during each iteration, half of the heads are randomly selected to be frozen. This implies that the parameters within these selected heads will remain unchanged and not undergo updates. Every head has an equal probability of being chosen. Before going to the next iteration, those frozen heads are unfrozen. The algorithm of the DiverseHead semi-supervised learning approach with the perturbation of dynamic freezing is expressed in the Algorithm 1.
Another form of perturbation involves the utilization of dropout layers. We introduce a dropout layer before the last convolutional layer for each head. The pseudocode of this method is expressed in Algorithm 2. In the training phase, specific components of the feature tensor are randomly assigned a value of zero with a dropout rate (determining the probability), using samples derived from a Bernoulli distribution. This dropout operation is employed to enhance the variability of the output predictions. In the conducted experiments, the dropout rate is set as a hyperparameter with a value of 0.3. Similar to the training procedure of the semi-supervised learning approach based on dynamic freezing, the dropout-based approach adheres to the same training pipeline outlined above, with the distinction that all heads remain unfrozen. To evaluate the efficacy of individual perturbation methods in conjunction with the proposed DiverseHead techniques, each method is applied independently within the proposed framework. The performance of each combination is assessed in Section V.
@@
[…]
        $\left\{head^{i}\right\}_{i=random}^{len(heads)/2}\leftarrow sample\left(\left\{head^{i}\right\}_{i=1}^{len(heads)}\right)$
        $Freeze(\left\{head^{i}\right\}_{i=random}^{len(heads)/2})$
        for $j\in\left\{1,...,len(heads)\right\}$ do
@@
        $r_{k}^{sum}=sum\left(\left\{r_{k}\right\}_{k=1}^{len(heads)}\right)$
        $\hat{y}_{k}^{sum}\leftarrow argmax(r_{k}^{sum})$
        $r_{k}^{main}\leftarrow sample\left(\left\{r_{k}^{1},r_{k}^{2},...,r_{k}^{len(heads)}\right\}\right)$
        $\mathcal{L}_{unsup}\leftarrow loss(r_{k}^{main},\hat{y}_{k}^{sum})$ based on (3)
        $\mathcal{L}\leftarrow\frac{1}{len(heads)}\sum_{j=1}^{len(heads)}\mathcal{L}_{sup}^{j}+\lambda\mathcal{L}_{unsup}$
@@
        $Unfreeze(\left\{head^{i}\right\}_{i=random}^{len(heads)/2})$
@@
[…]
        for $j\in\left\{1,...,len(heads)\right\}$ do
@@
        $r_{k}^{sum}=sum\left(\left\{r_{k}\right\}_{k=1}^{len(heads)}\right)$
        $\hat{y}_{k}^{sum}\leftarrow argmax(r_{k}^{sum})$
        $r_{k}^{main}\leftarrow sample\left(\left\{r_{k}^{1},r_{k}^{2},...,r_{k}^{len(heads)}\right\}\right)$
        $\mathcal{L}_{unsup}\leftarrow loss(r_{k}^{main},\hat{y}_{k}^{sum})$ based on (3)
        $\mathcal{L}\leftarrow\frac{1}{len(heads)}\sum_{j=1}^{len(heads)}\mathcal{L}_{sup}^{j}+\lambda\mathcal{L}_{unsup}$
@@
III Cross-model Supervision (DiverseModel)
@@
The proposed DiverseModel differs from CPS, exploring different networks in parallel to generate comprehensive pseudo labels. Since different networks pay different and complementary attention to the same input, this offers the basis that they are able to benefit from each other. DiverseModel structure includes three distinct networks which are PSPNet [zhao2017pyramid], UNet [ronneberger2015u], SegNet [badrinarayanan2017segnet], respectively. Algorithm 3 presents the pseudocode of the DiverseModel method.
@@
$\mathcal{L}_{sup}=\sum_{n=1}^{3}\frac{1}{W\times H}\sum_{k=1}^{W\times H}\left(\ell_{ce}\left(p_{k}^{n},y_{k}\right)\right),$ (5)
where $p_{k}^{n}$ represents the prediction of $k\leavevmode\nobreak\ th$ pixel from $n\leavevmode\nobreak\ th$ network.
In addition, unlabelled data is used to generate pseudo labels, which are then exploited for cross-supervision to inform each network. Different from the version presented in [ma2023confidence], all loss calculations in this work solely focus on cross-entropy loss to avoid the effects resulting from different types of loss functions on performance. The predictions obtained by each network are denoted as $\{p^{1},p^{2},p^{3}\}$, which are used for generating pseudo labels $\{\hat{r}^{1},\hat{r}^{2},\hat{r}^{3}\}$ through $argmax$ operation. For instance, the cross pseudo supervision loss $\mathcal{L}_{unsup}^{12}$ between the prediction $p^{1}$ from first network and the pseudo label $r^{2}$ generated by the second network is defined as:
$\mathcal{L}_{unsup}^{12}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\left(\ell_{ce}\left(p_{k}^{1},\hat{r}_{k}^{2}\right)\right).$ (6)
@@
[…]
@@
The cross-pseudo supervision among three networks creates six losses by using the same way. The unsupervised loss $\mathcal{L}_{unsup}$ is the result of linearly adding the six individual losses, as shown below
$\begin{array}[]{r}\mathcal{L}_{unsup}=\mathcal{L}_{unsup}^{12}+\mathcal{L}_{unsup}^{13}+\mathcal{L}_{unsup}^{21}\\ +\mathcal{L}_{unsup}^{23}+\mathcal{L}_{unsup}^{31}+\mathcal{L}_{unsup}^{32}.\end{array}$ (7)
@@ IV Dataset Description
We employed diverse datasets to assess both the proposed techniques and state-of-the-art methods, specifically including (1) the ISPRS Potsdam dataset [rottensteiner2012isprs], (2) the DFC2020 dataset [robinson2021global], (3) the RoadNet dataset [liu2018roadnet], and (4) the Massachusetts Buildings dataset [MnihThesis]. In the sequel, we share the details of each datasets utilised in this paper.
@@
1) ISPRS Potsdam Semantic Labeling dataset is an open-access benchmark dataset provided by the International Society for Photogrammetry and Remote Sensing (ISPRS). The ground sampling distance of the two modalities of the true orthophoto (TOP) and the DSM is 5 cm. This dataset was annotated manually into six land cover classes, which are impervious surfaces, buildings, low vegetation, trees, cars, and clutter/background. This dataset provides 38 multi-source patches (all with the size of $6000\times 6000$ pixels), which contain infrared (IR), red, green and blue bands orthorectified optical images with corresponding digital surface models (DSM). For computation purposes, we partitioned all these data tiles into $512\times 512$ patches, resulting in 3456, 201, and 1815 samples for training, validation, and test sets, respectively. We randomly select 1/4 of training data as labelled data and use the remaining 3/4 as unlabelled data for semi-supervised learning.
@@
2) DFC2020 is the 2022 IEEE GRSS Data Fusion Contest dataset which is based on the SEN12MS dataset [schmitt2019sen12ms]. It provides Sentinel-1 SAR imagery, Sentinel-2 multispectral imagery, and corresponding land cover maps across seven areas namely Mexico City, Kippa-Ring, Khabarovsk, Black Forest, Mumbai, Cape Town, Bandar Anzali between 2016 and 2017. The size of all patches is $256\times 256$ pixels. There are 6112, 986 and 5127 images for training, validation and test sets, respectively. In this paper, we utilize only one-fifth of the labelled data for training, while the remaining four-fifths are employed as unlabeled data. The fine-grained International Geosphere-Biosphere Program (IGBP) classification scheme in SEN12MS was aggregated to ten coarser-grained classes, which are forest (11.3%), shrubland (6.9%), savanna (23.6%), grassland (16.8%), wetlands (1.1%), croplands (17.9%), urban/built-up (10.6%), snow/ice (0.0%), barren (5.2%), and water (6.5%).
@@
3) RoadNet is a benchmark dataset for road detection with 0.21-m spatial resolution per pixel covering 21 typical urban regions about 8 $km^{2}$ in Ottawa, Canada. The annotation data includes road surfaces, road edges, and road centerlines for each image. we only used road surfaces for the segmentation task. the number of samples for training, validation and testing is 410, 45 and 387, respectively. In the semi-supervised learning approaches, only 1/4 of the annotated training data is used, while the remaining 3/4 of the training data are employed as unlabeled data for training purposes.
@@
4) Massachusetts Buildings Dataset predominantly encompasses urban and suburban regions, encompassing structures of varying scales. This includes diverse buildings such as individual houses and garages, all of which are included in the labels. The dataset consists of 151 aerial images capturing the Boston area covering roughly 340 square kilometres, with a resolution of $1\leavevmode\nobreak\ m^{2}$ per pixel. The dataset was divided into subsets as follows: 137 images for training, 10 images for testing, and 4 images for validation. Each subset consists of images sized $1500\times 1500$ pixels. One-quarter of the annotated training images and the remaining three-quarters of the unlabelled training data are utilized in semi-supervised learning.
V experiment and results analysis
V-A Implementation Details

### v2

@@ 3 DiverseHead (Cross-head Supervision)
Figure 2: DiverseHead: an online semi-supervised learning approach. This figure applies the dynamic freezing strategy: slide windows of the freezer randomly select a certain number of heads to freeze the parameter of heads (not updated by backpropagation). Additionally, during every iteration, all heads undergo supervision through a supervised loss, yet only a single head is randomly chosen to be updated by unsupervised loss.
Figure 3: The Proposed Voting Module: a voting mechanism for the pseudo label creation. In the unsupervised part, the voting module combines the mean output of multiple heads (mean voting) and individual pseudo labels (max voting) to generate more efficient pseudo labels. Argmax returns the indices of the maximum values of the prediction along the class dimension. The dashed arrow serves as an illustration of a pixel voting for its classification in a segmentation map.
@@
This section introduces the details of the proposed DiverseHead semi-supervised learning method which is proposed to create various predictions by multiple heads illustrated in Figure 2, specifically the number of convolutional layers for each head is 2. The source of perturbation is from two strategies, called dynamic freezing and dropout which are shown in Figure 4 and explained in detail in Section 3.1. The proposed DiverseHead does not distinguish between the main or auxiliary decoder, which differs from CCT. During every training iteration, all heads undergo supervision through supervised losses, yet only a single head is randomly chosen to be updated by an unsupervised loss. In addition, the proposed method combines self-training and consistency regularization, which uses their predictions to supervise itself and forces all perturbed head to give a consistent output. It makes an efficient voting module shown in Figure 3 to help calculate unsupervised loss by using optimized pseudo labels. To specifically explain the proposed semi-supervised framework, given both a labelled dataset $\mathcal{B}^{l}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$ containing $M$ images and an unlabelled data set $\mathcal{B}^{u}=\left\{u_{i}\right\}_{i=1}^{N}$ with $N$ images, the network $Q$ is constructed with multiple heads denoted as $\left\{head^{i}\right\}_{i=1}^{L}$, where L is the number of heads. Each of these heads is initialized differently. The expectation of the proposed semi-supervised learning architecture is to gain the trained network $Q$ leveraging these labelled and unlabelled data.
@@
The output of each head refers to a probability-based prediction for all classifications corresponding to its input. When working with labelled data, the supervised loss $\mathcal{L}_{sup}^{j}$ between the $k^{\prime}th$ ground truth $y_{k}$ and its corresponding prediction $p_{k}^{j}$ for $j^{\prime}th$ head is defined by using the standard cross-entropy loss function $\ell_{ce}$:
$\mathcal{L}_{sup}^{j}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\ell_{ce}\left(p_{k}^{j},y_{k}\right),$ (1)
@@
where $W$ and $H$ refer to the weight and height of input images. The final supervised loss is determined by the mean of the losses across all heads
$\mathcal{L}_{sup}=\frac{1}{L}\sum_{j=1}^{L}\mathcal{L}_{sup}^{j}$ (2)
where L is the number of heads.
In each iteration, we also get prediction $r_{k}^{j}$ corresponding to unlabelled data $u_{k}^{j}$ from the $j^{\prime}th$ head. Inspired by ensemble learning, a voting module is proposed to get high-precision pseudo labels. There are two voting mechanisms in the proposed voting module, called mean voting and max voting, respectively. The former aggregates the predictions from all heads to create a combined prediction $\hat{r}_{k}^{mean}$. Subsequently, this combined prediction is used to calculate the mean pseudo labels $\hat{y}_{k}^{mean}$ through an $argmax$ operation, which returns the indices of the maximum values of the prediction along the class dimension. Apart from the mean voting module, the individual pseudo label is generated from the output of each head by the $argmax$ function. The latter regards all pseudo labels as voters, in which the mean pseudo label contributes $\varphi$ weight and each individual pseudo label contributes unit weight. $\varphi$ is a learnable parameter and its value changes depending on the dataset and training processing. The max voting module returns the class number that gets the most votes for each pixel. After max voting, an optimal pseudo label $\hat{y}_{k}^{final}$ is created to calculate unsupervised loss by using the cross-entropy loss function:
@@
$\mathcal{L}_{unsup}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\ell_{ce}\left(r_{k}^{main},\hat{y}_{k}^{final}\right),$ (3)
where $r_{k}^{main}$ is the prediction from the selected head. There is a random selector to pick only one head as the main branch to be supervised by minimizing the unsupervised loss.
@@
(a) Dynamic Freezing
(b) Dropout
Figure 4: The proposed two perturbation methods: dropout and dynamic freezing. In (a), the dynamic selector is designed to amplify the parameter diversity of multiple heads. It randomly chooses a set number of heads not to be updated by backpropagation during each training iteration, thereby enhancing the variability of parameters.
3.1 Perturbation Methods
@@
  $L=length(heads)$
  for $\left\{\left(x_{k},y_{k}\right),u_{k}\right\}_{k=1}^{P}\in\left\{\textit{cycle}\left(\mathcal{B}^{l}\right),\mathcal{B}^{u}\right\}$ do
        $\mathcal{R}=Randint(0,L,\frac{1}{2}L)$
        $Freeze(\left\{head^{i}\right\}_{i\in\mathcal{R}})$
        for $j\in\left\{1,...,L\right\}$ do
@@
            $\hat{y}_{k}^{j}\leftarrow argmax(r_{k}^{j})$
        $r_{k}^{mean}=sum\left(\left\{r_{k}^{j}\right\}_{j=1}^{L}\right)$
        $\hat{y}_{k}^{mean}\leftarrow argmax(r_{k}^{mean})$
        $\hat{y}_{k}^{final}\leftarrow voting(\left\{\hat{y}_{k}^{j}\right\}_{j=1}^{L},\hat{y}_{k}^{mean})$
        $r_{k}^{main}\leftarrow sample\left(\left\{r_{k}^{1},r_{k}^{2},...,r_{k}^{L}\right\}\right)$
        $\mathcal{L}_{unsup}\leftarrow loss(r_{k}^{main},\hat{y}_{k}^{final})$ based on (3)
        $\mathcal{L}\leftarrow\frac{1}{L}\sum_{j=1}^{L}\mathcal{L}_{sup}^{j}+\lambda\mathcal{L}_{unsup}$
@@
        $\mbox{\it Unfreeze}(\left\{head^{i}\right\}_{i=R})$
@@
Based on the framework of DiverseHead, we propose a parameter perturbation method called dynamic freezing as shown in Figure 4 (a). The pseudocode of the algorithm for DiverseHead with the perturbation of dynamic freezing is given in Algorithm 1. Specifically, we use DeepLabv3+ with a ResNet-50 backbone as the segmentation network in our framework, ensuring consistency with the methods compared in this paper. The backbone is pretrained on ImageNet, and other parameters are initialised by Kaiming Initialization. Both labelled and unlabelled data are input to the framework, then the first step of dynamic freezing refers to a process where half of the heads are randomly selected to be frozen during each iteration. This implies that the parameters within these selected heads will remain unchanged and not undergo updates for the current iteration. Every head has an equal probability of being chosen. Following the cross-head supervision above, the segmentation model is updated by the supervised and unsupervised losses. Before going to the next iteration, those frozen heads are unfrozen.
@@
  $L=length(heads)$
  for $\left\{\left(x_{k},y_{k}\right),u_{k}\right\}_{k=1}^{P}\in\left\{\textit{cycle}\left(\mathcal{B}^{l}\right),\mathcal{B}^{u}\right\}$ do
        for $j\in\left\{1,...,L\right\}$ do
@@
            $\hat{y}_{k}^{j}\leftarrow argmax(r_{k}^{j})$
        $r_{k}^{sum}=sum\left(\left\{r_{k}^{j}\right\}_{j=1}^{L}\right)$
        $\hat{y}_{k}^{mean}\leftarrow argmax(r_{k}^{mean})$
        $\hat{y}_{k}^{final}\leftarrow voting(\left\{\hat{y}_{k}^{j}\right\}_{j=1}^{L},\hat{y}_{k}^{mean})$
        $r_{k}^{main}\leftarrow sample\left(\left\{r_{k}^{1},r_{k}^{2},...,r_{k}^{L}\right\}\right)$
        $\mathcal{L}_{unsup}\leftarrow loss(r_{k}^{main},\hat{y}_{k}^{final})$ based on (3)
        $\mathcal{L}\leftarrow\frac{1}{L}\sum_{j=1}^{L}\mathcal{L}_{sup}^{j}+\lambda\mathcal{L}_{unsup}$
@@
Another form of perturbation involves the use of dropout layers to the proposed DiverseHead structure to increase the diversity of features during training. We introduce a dropout layer after each convolutional block in each segmentation head of the network as shown in Figure 4 (b). The pseudocode of this method is given in Algorithm 2. As both dynamic freezing and dropout perturbation utilize the same semi-supervised learning framework, DiverseHead, the network initialization and supervised loss calculation procedures stay the same. Using dropout in DiverseHead, specific components of the weights in the heads of the network are randomly assigned a value of zero with a dropout rate (determining the probability), using samples derived from a Bernoulli distribution. This dropout operation is employed to enhance the variability of the output predictions. In the conducted experiments, the dropout rate is set as a hyperparameter with a value of 0.3. Although the dropout-based approach adheres to the same training pipeline with dynamic freezing outlined above, the distinction is that all heads remain unfrozen.
To evaluate the efficacy of individual perturbation methods in conjunction with the proposed DiverseHead techniques, each method is applied independently within the proposed framework. The performance of each combination is assessed in Section 6. Considering various datasets may require differing levels of diversity in pseudo labels within the proposed DiverseHead framework, adjustments can be made by altering the number of frozen heads and dropout rates in the methods of dynamic freezing and dropout, respectively.
4 Cross-model Supervision (DiverseModel)
@@
The proposed DiverseModel differs from CPS, exploring different networks in parallel to generate comprehensive pseudo labels. As shown in Figure 5, the DiverseModel structure includes three distinct networks, which can be various semantic segmentation networks. In this paper, we choose three widely used segmentation networks for experiments, which are PSPNet (Zhao et al., 2017), UNet (Ronneberger et al., 2015), SegNet (Badrinarayanan et al., 2017). Algorithm 3 presents the pseudocode of the DiverseModel method. Since different networks pay different and complementary attention to the same input, this offers the basis that they are able to benefit from each other. In order to provide evidence for this claim, we executed the Gradient-weighted Class Activation Mapping (Grad-CAM) (Selvaraju et al., 2020) technique for every network employed within the framework of the DiverseModel architecture. Grad-CAM visualizes the areas of an image that are important to the model predictions from each network. Figure 6 depicts an example grad-CAM analysis for the BUILDING class in the Potsdam data set. Examining Figure 6, we can see that different networks pay different and complementary attention to the same input, and the pseudo labels from the DiverseModel prediction show the highest quality.
Figure 6: The upper section displays Grad-CAM outputs from individual networks within the DiverseModel architecture using the Potsdam dataset. The lower section showcases both the ground truth and pseudo labels generated through predictions from each network. In the lower right corner, the pseudo label from DiverseModel is presented.
@@
$\mathcal{L}_{sup}=\frac{1}{3}\sum_{n=1}^{3}\frac{1}{W\times H}\sum_{k=1}^{W\times H}\ell_{ce}\left(p_{k}^{n},y_{k}\right),$ (5)
where $p_{k}^{n}$ represents the prediction of the $k\leavevmode\nobreak\ th$ pixel from the $n\leavevmode\nobreak\ th$ network.
In addition, unlabelled data is used to generate pseudo labels, which are then exploited for cross-supervision to inform each network. Different from the version presented in (Ma et al., 2023), all loss calculations in this work solely focus on cross-entropy loss to avoid the variations in performance resulting from different types of loss functions. The predictions obtained by each network are denoted as $\{p^{1},p^{2},p^{3}\}$, which are used for generating pseudo labels $\{\hat{r}^{1},\hat{r}^{2},\hat{r}^{3}\}$ through the $argmax$ operation. For instance, the cross pseudo supervision loss $\mathcal{L}_{unsup}^{12}$ between the prediction $p^{1}$ from the first network and the pseudo label $r^{2}$ generated by the second network is defined as:
$\mathcal{L}_{unsup}^{12}=\frac{1}{W\times H}\sum_{k=1}^{W\times H}\ell_{ce}\left(p_{k}^{1},\hat{r}_{k}^{2}\right).$ (6)
@@
  for $\left\{\left(x_{k},y_{k}\right),u_{k}\right\}_{k=1}^{P}\in\left\{\textit{cycle}\left(\mathcal{B}^{l}\right),\mathcal{B}^{u}\right\}$ do
@@
The cross-pseudo supervision among three networks creates six losses in the same way. The unsupervised loss $\mathcal{L}_{unsup}$ is the average of the six individual losses, as shown below
$\begin{array}[]{r}\mathcal{L}_{unsup}=\frac{1}{6}\left(\mathcal{L}_{unsup}^{12}+\mathcal{L}_{unsup}^{13}+\mathcal{L}_{unsup}^{21}+\mathcal{L}_{unsup}^{23}+\mathcal{L}_{unsup}^{31}+\mathcal{L}_{unsup}^{32}\right).\end{array}$ (7)
@@ 5 Dataset Description
We employed diverse remote sensing datasets to assess both the proposed techniques and state-of-the-art methods, specifically including (1) the ISPRS Potsdam dataset (Rottensteiner et al., 2012), (2) the DFC2020 dataset (Robinson et al., 2021), (3) the RoadNet dataset (Liu et al., 2018), and (4) the Massachusetts Buildings dataset (Mnih, 2013). In the sequel, we share the details of each dataset utilised in this paper.
@@
1) ISPRS Potsdam Semantic Labeling dataset is an open-access benchmark dataset provided by the International Society for Photogrammetry and Remote Sensing (ISPRS). The ground sampling distance of the two modalities of the true orthophoto (TOP) and the DSM is 5 cm. This dataset was annotated manually into six land cover classes, which are impervious surfaces, buildings, low vegetation, trees, cars, and clutter/background. This dataset provides 38 patches (all of the size $6000\times 6000$ pixels), which contain infrared (IR), red, green and blue bands orthorectified optical images with corresponding digital surface models (DSM). For computation purposes, we partitioned all these data tiles into $512\times 512$ patches, resulting in 3456, 201, and 1815 samples for training, validation, and test sets, respectively. We randomly select a quarter of the training data as labelled data and use the remaining three quarters as unlabelled data for semi-supervised learning.
@@
2) DFC2020 is the 2022 IEEE GRSS Data Fusion Contest dataset which is based on the SEN12MS dataset (Schmitt et al., 2019). It provides Sentinel-1 SAR imagery, Sentinel-2 multispectral imagery, and corresponding land cover maps across seven areas namely Mexico City, Kippa-Ring, Khabarovsk, Black Forest, Mumbai, Cape Town, Bandar Anzali between 2016 and 2017. The size of all patches is $256\times 256$ pixels. There are 6112, 986 and 5127 images for training, validation and test sets, respectively. In this paper, we utilize only one-fifth of the labelled data for training, while the remaining four-fifths are employed as unlabeled data. The fine-grained International Geosphere-Biosphere Program (IGBP) classification scheme in SEN12MS was aggregated into ten coarser-grained classes.
@@
3) RoadNet is a benchmark dataset for road detection with 0.21-m spatial resolution per pixel covering 21 typical urban regionsThe annotation data includes road surfaces, road edges, and road centerlines for each image. We only used road surfaces for the segmentation task. the number of samples for training, validation and testing is 410, 45 and 387, respectively. In the semi-supervised learning approaches, only a quarter of the annotated training data is used, while the remaining three-quarters of the training data are employed as unlabeled data for training purposes.
@@
4) Massachusetts Buildings Dataset predominantly encompasses urban and suburban regions, encompassing structures of varying scales. This includes diverse buildings such as individual houses and garages, all of which are included in the labels. The dataset consists of 151 aerial images capturing the Boston area covering roughly 340 square kilometres, with a resolution of $1\leavevmode\nobreak\ m^{2}$ per pixel. The dataset was divided into subsets as follows: 137 images for training, 10 images for testing, and 4 images for validation. Each subset consists of images sized $1500\times 1500$ pixels. A quarter of the annotated training images and the remaining three-quarters of the unlabelled training data are used in semi-supervised learning.
6 Experimental analysis
6.1 Implementation Details